OpenStudy (anonymous):

Select the best possible first step to solving the system by first eliminating the x variable. x – 3y = 12 2x + 5y = 2 Multiply the second equation by 2. Multiply the first equation by 2. Multiply the first equation by –2. Multiply the second equation by –2.

OpenStudy (anonymous):

Multiply the first equation by -2. Then add to the second.

OpenStudy (anonymous):

You should get: 11y = 14 y = 14/11, plug Y back into one of the equations and you will get your X value.

OpenStudy (anonymous):

x – 3y = 12 2x + 5y = 2 Multiply the first equation by –2. (and then add it to the second equation) -2x + 6y = -24 2x + 5y = 2 on adding them we get 11y = -22 y = -22/11 = -2 substitute y = -2 in first equation x - 3(-2) = 12 x + 6 = 12 x = 12 - 6 x = 6